# Minimalistic and Scalable Quantum Reservoir Computing Enhanced with Feedback

### v1

@@ Abstract
Quantum Reservoir Computing (QRC) leverages quantum systems to perform complex computational tasks with exceptional efficiency and reduced energy consumption. We introduce a minimalistic QRC framework utilizing only a few two-level atoms in a single-mode optical cavity, combined with continuous quantum measurements. To achieve high computational expressivity with minimal hardware, we include two critical elements: reservoir feedback and polynomial regression. Reservoir feedback modifies the reservoir’s dynamics without altering its hardware, while polynomial regression enhances output resolution by nonlinearly extending expressions. We evaluate QRC’s memory retention and nonlinear data processing through two tasks: predicting chaotic time-series data via the Mackey-Glass task and classifying sine-square waveforms. Our results demonstrate significant QRC performance with minimal reservoirs containing as few as five atoms, further enhanced by feedback mechanisms and polynomial regression. This framework fulfills QRC’s objectives to minimize hardware size and energy consumption, marking a significant advancement in integrating quantum physics with machine learning technology.
@@ I Introduction
Quantum reservoir computing (QRC) has emerged as a groundbreaking machine learning framework, capable of addressing intricate tasks with remarkable efficiency and minimal energy consumption Fujii2017 ; chen2019learning ; CNY20 ; Dudas2023 ; Govia2021 ; Bravo2022 ; Hulser2023 ; Nokkala2021 ; Pena2021 ; Xia2022 ; Mujal2023 ; Fry2023 ; Kalfus2022 ; Yasuda23 ; Beni2023 ; Markovic2019 ; Lin2020 ; kornjaca2024 . Some of these tasks fall into well-established categories in machine learning, such as time series analysis Fujii2017 ; Dudas2023 ; Govia2021 and computer vision kornjaca2024 , providing a clear benchmark for comparing the performance of QRC against classical reservoir computing (CRC) Tanaka2019 ; Appeltant2011 ; Pathak2018 ; Angelatos2021 ; Chen2022 ; Gauthier2021 . In addition, various QRC approaches have been employed to tackle tasks in the field of quantum research, such as recognizing quantum entanglements Ghosh2019 , measuring dispersive currents Angelatos2021 , and predicting molecular structures Domingo2022 , often referred to as quantum random kitchen sinks or quantum extreme learning machines Innocenti23 .
@@
Proposals for implementing quantum reservoirs have been advanced using various platforms, including coupled networks of qubits Fujii2017 ; chen2019learning ; CNY20 ; Domingo2022 ; Ghosh2021 ; Pena2021 ; Xia2022 ; Yasuda23 , fermions Ghosh2019 , harmonic oscillators Nokkala2021 , Kerr nonlinear oscillators Angelatos2021 , Rydberg atoms Bravo2022 , and optical pulses Beni2023 . Increasing the number of physical sites, such as qubits, oscillators, or atoms, within a quantum network can dramatically expand the Hilbert space by exponentially increasing the number of quantum basis states. Each of these basis states functions as a node within the quantum neural network, mirroring the role of a node in a classical neural network Fujii2017 or an optical neural network Wang2022 ; ma2023 . The exponential scaling of basis states plays a critical role in making the system’s dynamics nonlinear and complex. To ensure QRC performance, tens of physical sites are typically necessary for constructing a quantum reservoir kornjaca2024 . However, a significant challenge in scaling up a quantum reservoir is the considerable difficulty of establishing connections between each newly added site and all existing sites within a quantum network. To extract information from quantum reservoirs, the observable readouts have been proposed through various methods, including measuring probabilities on quantum basis states Dudas2023 , excitations of qubits Fujii2017 , occupations on lattice sites Ghosh2019 , and energies of oscillators Angelatos2021 . However, these conventional measurement techniques often require quantum tomography, which necessitates the complete destruction of the quantum reservoir to obtain readouts at each time step. This process demands a substantial number of repeated time evolutions, incurring a considerable overhead.
@@
In this paper, we present a minimalistic quantum reservoir made of up to five atoms and a single-mode cavity field, where the input is practically embedded in external coherent laser driving. The connections among atoms can be more conveniently constructed compared to prior proposals, since newly added atoms automatically couple to existing atoms through their mutual connections with the cavity field.
@@
In small hardware systems, maximizing computational expressivity is imperative. To achieve this, we adopt two main strategies: adjustable feedback and polynomial regression. First, we introduce a feedback mechanism that feeds observable readouts back into the reservoir as input. This approach is related to the feedback formalism in classical echo state networks (ESNs) Peter2023 and biological systems Maass2007 . This mechanism provides an excellent solution: it allows us to modify the reservoir’s overall dynamics and significantly enhance its computational expressivity without altering its internal hardware details. Second, we utilize polynomial combinations of readouts to enrich the complexity of the output, leading to another significant performance boost. The observable readouts in our scheme are practically obtained through a continuous quantum measurement, where the cavity field provides two readouts linked to two photonic quadratures, and each atom provides two readouts linked to two atomic spin channels Wiseman2010 ; BvHJ07 ; Wei2008 ; Ruskov2010 ; Hacohen2016 ; Ochoa2018 ; Fuchs2001 . By adopting a continuous measurement scheme that accounts for measurement back-action on the reservoir, we eliminate the need for the costly tomography measurements used in previous methods. Our proposed approach, implemented on minimalistic quantum hardware, combines a feedback mechanism, polynomial regression, and continuous measurement. Together, these elements aim to achieve highly efficient QRC with minimal energy consumption.
@@
We evaluate the QRC performance on two time-series tasks: the Mackey-Glass forecasting task Fujii2017 , which demands long-term memory for predicting the future trend of an input function while providing the fading memory property simultaneously, and the sine-square waveform classification task Dudas2023 ; Markovic2019 , which requires nonlinearity of the reservoir to capture sudden, high-frequency shifts in a linearly inseparable input dataset. The significant performance enhancement associated with the increased number of atoms, the feedback mechanism, and the polynomial regression are demonstrated by the diminishing gaps between the actual and target outputs in the two tasks.
@@ II.1 Quantum Reservoir Computing Setup
We consider a QRC framework with a feedback mechanism, as illustrated in Fig. 1. In this setup, input samples are encoded in a classical function $f_{k}$, where $k$ is the time-series index. The input is processed by the quantum reservoir, and the corresponding readouts $x_{kn}$ are acquired through continuous quantum measurement, with $n$ representing the index of the readout channel. The feedback mechanism is implemented by using a subset of the readouts, denoted as $x_{kn_{f}}$, to modify the input as
@@
where $n_{f}$ indexes the feedback channels. The readouts are then linked to the output to perform a time-series forecasting or classification task via polynomial regression, where the associated weights are denoted as $W_{n}$. The machine-learning parameters $V_{n_{f}}$ and $W_{n}$ are trained using global optimization and pseudoinverse regression, respectively, to optimize the QRC performance.
We employ a quantum optical system encompassing $N_{atom}$ two-level atoms inside a single-mode optical cavity Jaynes1963 ; Tavis1968 as the quantum reservoir. The time-independent part of the Hamiltonian is expressed as
[…]
where $c$ represents the photon annihilation operator, $\sigma_{i}=\left|g\right\rangle\left\langle e\right|_{i}$ denotes the lowering operator of the $i$-th atom with $\left|g\right\rangle$ ($\left|e\right\rangle$) representing the ground (excited) state, $\omega_{c}$ ($\omega_{i}$) describes the detuning between the coherent driving and the cavity (atomic) frequencies, and $g_{i}$ is the electric-dipole coupling strength between the $i$-th atom and the cavity mode. For QRC, it is essential to select either various detuning $\omega_{i}$ or various coupling strength $g_{i}$ to prompt the atoms to produce non-identical memories, thus enhancing their overall capability. This model can be practically implemented using cold atoms in an optical cavity Brennecke2007 ; Niemczyk2010 , where optical tweezers can be used to trap and measure individual atoms at different positions Kaufman2021 ; Ye2023 . Alternatively, quantum dots can also be utilized to induce random positioning, detuning, and coupling Schmidt2007 .
@@
The readouts $x_{kn}$ are determined through continuous measurement of quantum observables. The observables of the cavity field stem from the homodyne detection of two orthogonal quadratures Wiseman2010 ; BvHJ07 ; Nurdin14
@@
and the observables of the atomic spontaneous emission are associated with the Pauli operators Wiseman2001
@@
It has been demonstrated that these observables of the cavity and atoms can be simultaneously measured Wei2008 ; Ruskov2010 ; Hacohen2016 ; Ochoa2018 . By averaging over a large number of measurement trajectories, the readouts are correlated with the expectation values as $x_{k1}=\left\langle Q\right\rangle_{k}$, $x_{k2}=\left\langle P\right\rangle_{k}$, $x_{k3}=\left\langle\sigma_{x,1}\right\rangle_{k}$, $x_{k4}=\left\langle\sigma_{y,1}\right\rangle_{k}$, and so forth. This leads to $N_{readouts}=2N_{atom}+2$, where $N_{readouts}$ is the number of available readout channels and $N_{atom}$ is the number of atoms. The expectation values are calculated as $\left\langle Q\right\rangle_{k}={\rm Tr}\left[\rho\left(t_{k}\right)Q\right]$, for instance, where $\rho\left(t_{k}\right)$ represents the density operator at time $t_{k}$, which evolves according to the dynamics of the reservoir simulated by the quantum master equation. The dynamics of the quantum reservoir with continuous quantum measurement are elaborated in the "Methods" section.
To implement the feedback mechanism, a subset of readouts, denoted as $x_{kn_{f}}$, is selected to serve as feedback. These readouts are combined with the original input $f_{k}$ to generate the modified input $\widetilde{f}_{k}$, as described in Eq. (1). At time $t_{k}$, the modified input enters the quantum reservoir in terms of the coherent driving shown in Eq. (3). The evolution of the quantum reservoir from $t_{k}$ to $t_{k+1}$, simulated by the master equation, produces the readouts $x_{k+1,n_{f}}$ at time $t_{k+1}$, which are then used to modify the input $\widetilde{f}_{k+1}$ for the next feedback cycle.
@@
The goal of training is to optimize the parameters $V_{n_{f}}$ and $W_{n}$ to minimize the normalized root mean square error (NRMSE). We adopt various global optimization methods to train $V_{n_{f}}$ and the Moore-Penrose pseudoinverse method to train $W_{n}$. The definition of NRMSE and the specific training methods are detailed in the "Methods" section. During the testing phase, the trained weights $V_{n_{f}}$ and $W_{n}$ are treated as fixed, and the NRMSE is computed to evaluate the QRC performance.
@@ II.2 Mackey-Glass Task
The Mackey-Glass task serves as a test for long-term memory, demanding the reservoir to retain past information from the input function to forecast its future behavior. The input function is generated from the Mackey-Glass equation
@@
where the parameters $\beta=0.2$, $\gamma=0.1$, and $\tau_{M}=17$ are commonly accepted standard values in the chaotic regime Fujii2017 ; Dudas2023 ; Hulser2023 . We apply a buffer by discarding the first $1000$ time units in Eq. (9), thus considering the discretized function $f_{k}=f\left(t+1000\right)$ as the original input, with discretized time $t=k\Delta t$, as depicted in Fig. 2(a). The time series is sampled with a time step of $\Delta t=1$. The actual input, $\widetilde{f}_{k}$, is modified by the feedback system in each time step, as described by Eq. (1). The target output, $\bar{y}_{k}$, predicting the future of the input function with a time $Delay$, is constructed as $\bar{y}_{k}=f_{k+Delay}$. The readouts of observables illustrated in Fig. 2(c)-(f) show that the reservoir exhibits discernible responses to various input waveforms, which are exemplified by the green shadows corresponding to two different input waveforms. The time series is divided into three intervals for the purposes of memory fading, training and testing. The memory fading phase guarantees that during training and testing, the readouts depend solely on the input function rather than the initial state of the master equation.
@@
Figure 3: Performance testing for the Mackey-Glass task. (a)(b) The tested NRMSE plotted against the number of atoms (or neurons), using all available readouts from the cavity field and all atoms, with no feedback (black), $2$ feedbacks via $x_{k1}$ and $x_{k2}$ channels (blue), or $4$ feedbacks via $x_{k1}$ to $x_{k4}$ channels (green), where regular linear regression (solid lines) or polynomial regression (dashed lines) is applied. In panel (a), the coupling strength $g_{i}=30$ is fixed for all atoms, and the detunings $\omega_{i}$ vary: $\omega_{i}=20$ for one atom, $\omega_{i}=[0,40]$ for two atoms, $\omega_{i}=[0,20,40]$ for three atoms, $\omega_{i}=[0,10,30,40]$ for four atoms, and $\omega_{i}=[0,10,20,30,40]$ for five atoms. In panel (b), $\omega_{i}=20$ is fixed for all atoms, and $g_{i}$ varies: $g_{i}=30$ for one atom, $g_{i}=[10,50]$ for two atoms, $g_{i}=[10,30,50]$ for three atoms, $g_{i}=[10,20,40,50]$ for four atoms, and $g_{i}=[10,20,30,40,50]$ for five atoms. In panels (a)(b), classical reservoir computing (CRC) (red solid line) uses all available readouts from all neurons in classical echo state networks (ESN) (detailed in the "Supplemental Information"). (c) The tested NRMSE as a function of the number of unmeasured atoms (for QRC) or unmeasured neurons (for CRC), using only the $4$ readout channels (taking 4 measurements) from the cavity field and a particular atom with $g=30$ and $\omega=20$, where the total number of atoms is increased by fixing either $g_{i}$ (solid lines) or $\omega_{i}$ (dashed-dotted lines). (d)(e) Actual (red) and target (blue) outputs from QRC utilizing one atom with no feedback with linear regression, and three atoms with $4$ feedbacks with polynomial regression, corresponding to the points marked by letters "d" and "e" in panel (b), respectively. (f)(g) The influence of different selections for the $2$ feedback channels out of the $6$ readout channels in two-atom QRC (black lines), where ($n$,$m$) denotes the feedback channels $x_{kn}$ and $x_{km}$ with $n<m$. The results from one-atom and three-atom QRC using feedback channels $x_{k1}$ and $x_{k2}$ (orange lines) are plotted for reference. Panel (f) fixes $g_{i}$ for all atoms, while panel (g) fixes $\omega_{i}$. (h) Comparison of three methods for training feedback parameters $V_{n_{f}}$: differential evolution, brute force, and brute force plus Nelder Mead, under varying numbers of feedbacks for one-atom QRC with $g_{1}=30$ and $\omega_{1}=20$ using linear regression. Parameters: $Delay=20$, $\kappa=10$, $\omega_{c}=40$, $\epsilon=20$.
The scale of QRC is determined by four factors: (i) the number of basis states spanning the Hilbert space, which scales as $N_{c}2^{N_{atom}}$, where $N_{c}$ is the number of involved photon Fock states and $N_{atom}$ is the total number of "measured" and "unmeasured" atoms; (ii) the number of readouts, analogous to the number of neurons in CRC, which is determined by the number of "measured" atoms, with a maximum availability of $22N_{atom}+$ 2; (iii) the number of feedbacks, which is decided by the number of readouts that are used to modify the input through the feedback mechanism; and (iv) the implementation of polynomial regression, which incorporates both linear and quadratic terms of the readouts.
@@
The performance enhancement on account of the scalability of QRC is illustrated in Fig. 3. Increasing the number of atoms boosts the performance by expanding both the size of the Hilbert space and the number of available readouts. Figures 3(a)(b) show the results when the number of atoms is increased from $1$ to $5$, utilizing all available readout channels from the cavity field and all the atoms, with $g_{i}$ and $\omega_{i}$ fixed, respectively. This leads to a corresponding increase in the number of readouts (neurons) from $4$ to $12$. The remarkable performance enhancement associated with the feedback mechanism is depicted by the blue ($2$ feedbacks) and green ($4$ feedbacks) lines in Figs. 3(a)(b), while the results from polynomial regression are shown by the dashed lines. The feedback parameters $V_{n_{f}}$ have been trained using three global optimization methods, with the test results compared in Fig. 3(h) and details provided in the "Methods" section. Additionally, the NRMSE in the space of feedback parameters $V_{n_{f}}$ is illustrated in the "Supplemental Information", where high-frequency oscillations in the readouts, related to a limit-cycle-like behavior, are observed in certain parameter regions. A comparison between two extreme cases, one-atom QRC without feedback finished by linear regression and three-atom QRC with 4 feedbacks finished by polynomial regression, is illustrated in Figs. 3(d)(e). This comparison demonstrates a substantial performance improvement resulting from the combination of all the scalable factors.

### v2

@@ Abstract
Abstract
Quantum Reservoir Computing (QRC) leverages quantum systems to perform complex computational tasks with exceptional efficiency and reduced energy consumption. We introduce a minimalistic QRC framework utilizing as few as five atoms in a single-mode optical cavity, combined with continuous quantum measurement. The system is conveniently scalable, as newly added atoms naturally couple with existing ones via the shared cavity field. To achieve high computational expressivity with a minimal reservoir, we include two critical elements: reservoir feedback and polynomial regression. Reservoir feedback modifies the reservoir’s dynamics without altering its internal quantum hardware, while polynomial regression nonlinearly enhances output resolution. We demonstrate significant QRC performance in memory retention and nonlinear data processing through two tasks: predicting chaotic time-series data via the Mackey-Glass task and classifying sine-square waveforms. This framework fulfills QRC’s objectives to minimize hardware size and energy consumption, marking a significant advancement in integrating quantum physics with machine learning technology.
@@ I Introduction
Quantum reservoir computing (QRC) has emerged as a groundbreaking machine learning framework, capable of addressing intricate tasks with remarkable efficiency and minimal energy consumption [Fujii2017, chen2019learning, CNY20, Dudas2023, Govia2021, Bravo2022, Hulser2023, Nokkala2021, Pena2021, Xia2022, Mujal2023, Fry2023, Kalfus2022, Yasuda23, Beni2023, Markovic2019, Lin2020, kornjaca2024]. Some of these tasks fall into well-established categories in machine learning, such as time series analysis [Fujii2017, Dudas2023, Govia2021] and computer vision [kornjaca2024], providing a clear benchmark for comparing the performance of QRC against classical reservoir computing (CRC) [Tanaka2019, Appeltant2011, Pathak2018, Angelatos2021, Chen2022, Gauthier2021]. In addition, various QRC approaches and its static variants that process vectors rather than time series data, called quantum random kitchen sinks or quantum extreme learning machines [Innocenti23], have been employed to tackle tasks in the field of quantum research, such as recognizing quantum entanglements [Ghosh2019], measuring dispersive currents [Angelatos2021], and predicting molecular structures [Domingo2022].
@@
Proposals for implementing quantum reservoirs have been advanced using various platforms, including coupled networks of qubits [Fujii2017, chen2019learning, CNY20, Domingo2022, Ghosh2021, Pena2021, Xia2022, Yasuda23, McMahon2024], fermions [Ghosh2019], harmonic oscillators [Nokkala2021], Kerr nonlinear oscillators [Angelatos2021], Rydberg atoms [Bravo2022], microparticles [Cichos2024], chiral magnets [Kurebayashi20231, Kurebayashi20232], optical pulses [Beni2023], artificial spin systems [Kurebayashi2022], chemical reaction networks [Baltussen2024], and phonon and magnon modes [Scherbakov2023]. Increasing the number of physical sites, such as qubits, oscillators, or atoms, within a quantum network can dramatically expand the Hilbert space by exponentially increasing the number of quantum basis states. Each of these basis states functions as a node within the quantum neural network, mirroring the role of a node in a classical neural network [Fujii2017] or an optical neural network [Wang2022, ma2023]. The exponential scaling of basis states plays a critical role in making the system’s dynamics nonlinear and complex. To ensure QRC performance, tens of physical sites are typically necessary for constructing a quantum reservoir [kornjaca2024]. However, a significant challenge in scaling up a quantum reservoir is the considerable difficulty of establishing connections between each newly added site and all existing sites within a quantum network.
To extract information from quantum reservoirs, the observable readouts have been proposed through various methods, including measuring probabilities on quantum basis states [Dudas2023], excitations of qubits [Fujii2017], occupations on lattice sites [Ghosh2019], and energies of oscillators [Angelatos2021]. However, these conventional measurement techniques often require quantum tomography, which necessitates the complete destruction of the quantum reservoir to obtain readouts at each time step. This process demands a substantial number of repeated time evolutions, incurring a considerable overhead. To overcome this time complexity issue, several online protocols for QRC have been proposed. These include: online measurement based on partial readout and subsequent resetting of readout qubits to preserve memory of past inputs [CNY20, Yasuda23, Hu2024]; weak and projective measurements to extract information accurately without hindering memory [Mujal2023]; a feedback-driven QRC framework that feeds measurement outcomes back into the reservoir to restore memory of prior inputs [Kobayashi2024]; and the application of singular value decomposition and data-filtering techniques to optimize the training of the finite-sampled QRC framework [Ahmed2025].
@@
In this paper, we present a minimalistic quantum reservoir made of up to five atoms and a single-mode cavity field, where the input is practically embedded in external coherent laser driving. The connections among atoms can be more conveniently constructed compared to prior proposals, since newly added atoms automatically couple to existing atoms through their mutual connections with the cavity field, ensuring the practical scalablilty of the system.
@@
In small hardware systems, maximizing computational expressivity is imperative. To achieve this, we adopt two main strategies: adjustable feedback and polynomial regression. First, we introduce a feedback mechanism that feeds observable readouts back into the reservoir as input. This approach is related to the feedback formalism in classical echo state networks (ESNs) [Peter2023] and biological systems [Maass2007]. This mechanism provides an excellent solution: it allows us to modify the reservoir’s overall dynamics and significantly enhance its computational expressivity without altering its internal hardware details. Second, we utilize polynomial combinations of readouts to enrich the complexity of the output, leading to another significant performance boost. The observable readouts in our scheme are practically obtained through a continuous quantum measurement, where the cavity field provides two readouts linked to two photonic quadratures, and each atom provides two readouts linked to two atomic spin channels [Wiseman2010, BvHJ07, Wei2008, Ruskov2010, Hacohen2016, Ochoa2018, Fuchs2001]. By adopting a continuous measurement scheme that accounts for measurement back-action on the reservoir, we eliminate the need for the costly tomography measurements used in previous methods. Our proposed approach, implemented on minimalistic quantum hardware, combines a feedback mechanism, polynomial regression, and continuous measurement. Together, these elements aim to achieve highly efficient QRC with minimal energy consumption.
@@
We evaluate the QRC performance on two time-series tasks: the Mackey-Glass forecasting task [Fujii2017], which demands long-term memory for predicting the future trend of an input function while providing the fading memory property simultaneously, and the sine-square waveform classification task [Dudas2023, Markovic2019], which requires nonlinearity of the reservoir to capture sudden, high-frequency shifts in a linearly inseparable input dataset. The significant performance enhancement associated with the increased number of atoms, the feedback mechanism, and the polynomial regression are demonstrated by the diminishing gaps between the actual and target outputs in the two tasks.
@@ II.1 Quantum Reservoir Computing Setup
We consider a QRC framework with a feedback mechanism, as illustrated in Fig. 1. In this setup, input samples are encoded in a real-valued time series $f_{k}$, where the integer $k$ denotes the discrete time index. The input is processed by the quantum reservoir, and the corresponding readouts $x_{kn}$ are acquired through continuous quantum measurement, with $n$ representing the index of the readout channel. The feedback mechanism is implemented by using a subset of the readouts, denoted as $x_{kn_{f}}$, to modify the input as
@@
where $n_{f}$ indexes the feedback channels. The readouts are then linked to the output to perform a time-series forecasting or classification task via polynomial regression, where the associated weights are denoted as $W_{n}$. The machine-learning parameters $V_{n_{f}}$ and $W_{n}$ are trained using global optimization and pseudoinverse regression, respectively, to optimize the QRC performance. Note that the framework of Eq. (1) represents a general feedback formalism applicable to any quantum reservoir in which the input is modified by its own measurement outcomes. In the quantum optical reservoir considered below, $\widetilde{f}_{k}$ is implemented via coherent driving of a cavity, while $x_{kn_{f}}$ corresponds to continuous measurement of observable expectation values.
We employ a quantum optical system encompassing $N_{atom}$ two-level atoms inside a single-mode optical cavity [Jaynes1963, Tavis1968] as the quantum reservoir. The time-independent part of the Hamiltonian is expressed as
$H_{0}=\omega_{c}c^{\dagger}c+\underset{i}{\sum}\omega_{i}\sigma_{i}^{\dagger}\sigma_{i}+\underset{i}{\sum}g_{i}\left(c^{\dagger}\sigma_{i}+c\sigma_{i}^{\dagger}\right),$ (2)
where $c$ represents the photon annihilation operator, $\sigma_{i}=\left|g\right\rangle\left\langle e\right|_{i}$ denotes the lowering operator of the $i$-th atom with $\left|g\right\rangle$ ($\left|e\right\rangle$) representing the ground (excited) state, $\omega_{c}$ ($\omega_{i}$) describes the detuning between the coherent driving and the cavity (atomic) frequencies, and $g_{i}$ is the electric-dipole coupling strength between the $i$-th atom and the cavity mode. For QRC, it is essential to select either various detuning $\omega_{i}$ or various coupling strength $g_{i}$ to prompt the atoms to produce non-identical memories, thus enhancing their overall capability. This model can be practically implemented using cold atoms in an optical cavity [Brennecke2007, Niemczyk2010], where optical tweezers can be used to trap and measure individual atoms at different positions [Kaufman2021, Ye2023]. Alternatively, quantum dots can also be utilized to induce random positioning, detuning, and coupling [Schmidt2007].
@@
The readouts $x_{kn}$ are determined through continuous measurement of quantum observables. The observables of the cavity field stem from the homodyne detection of two orthogonal quadratures [Wiseman2010, BvHJ07, Nurdin14]
@@
and the observables of the atomic spontaneous emission are associated with the Pauli operators [Wiseman2001]
@@
It has been demonstrated that these observables of the cavity and atoms can be simultaneously measured [Wei2008, Ruskov2010, Hacohen2016, Ochoa2018]. By averaging over a large number of measurement trajectories, the readouts are correlated with the expectation values as $x_{k1}=\left\langle Q\right\rangle_{k}$, $x_{k2}=\left\langle P\right\rangle_{k}$, $x_{k3}=\left\langle\sigma_{x,1}\right\rangle_{k}$, $x_{k4}=\left\langle\sigma_{y,1}\right\rangle_{k}$, and so forth. This leads to $N_{readouts}=2N_{atom}+2$, where $N_{readouts}$ is the number of available readout channels and $N_{atom}$ is the number of atoms. The expectation values are calculated as $\left\langle Q\right\rangle_{k}={\rm Tr}\left[\rho\left(t_{k}\right)Q\right]$, for instance, where $\rho\left(t_{k}\right)$ represents the density operator at time $t_{k}$, which evolves according to the dynamics of the reservoir simulated by the quantum master equation. The dynamics of the quantum reservoir with continuous quantum measurement are elaborated in the “Methods” section.
To implement the feedback mechanism, a subset of readouts, denoted as $x_{kn_{f}}$, is selected to serve as feedback. These readouts are combined with the original input $f_{k}$ to generate the modified input $\widetilde{f}_{k}$, as described in Eq. (1). At time $t_{k}$, the modified input enters the quantum reservoir in terms of the coherent driving shown in Eq. (3). The evolution of the quantum reservoir from $t_{k}$ to $t_{k+1}$, simulated by the master equation, produces the readouts $x_{k+1,n_{f}}$ at time $t_{k+1}$, which are then used to modify the input $\widetilde{f}_{k+1}$ for the next feedback cycle. This reservoir feedback scheme can be viewed as a partial hidden state of a recurrent-structured reservoir using a fanned-out readout. However, compared to those in classical recurrent structures, our readouts, $x_{kn_{f}}$, are generated by a nonlinear quantum system with a large number of hidden basis states. This ensures significant QRC performance even with a small number of readouts since our simple feedback captures the complex dynamics occurring in a high-dimensional Hilbert space. Moreover, it is only partial feedback since not all of the expectation values of basis operators that span the space of all observables of the quantum system are fedback.
In the “Methods” section, we present an online protocol for the feedback mechanism based on continuous quantum measurement. This approach eliminates the need for multiple copies of the quantum system and repeated measurements to estimate the expectation values of observables at each time step. By enabling multiple continuous measurement trajectories on a single quantum system, the protocol avoids the use of destructive measurements.
@@
The goal of training is to optimize the parameters $V_{n_{f}}$ and $W_{n}$ to minimize the normalized root mean square error (NRMSE) defined by
${\rm NRMSE}=\frac{1}{\bar{y}_{max}-\bar{y}_{min}}\sqrt{\frac{\underset{k}{\sum}\left(y_{k}-\bar{y}_{k}\right)^{2}}{L}},$ (9)
where $L$ is the number of time steps in the training period and $\bar{y}_{k}$ represents the target output that captures some key features of the input. We adopt various global optimization methods to train $V_{n_{f}}$ and the Moore-Penrose pseudoinverse method to train $W_{n}$. The specific training methods are detailed in the “Methods” section. During the testing phase, the trained weights $V_{n_{f}}$ and $W_{n}$ are treated as fixed, and the NRMSE is computed to evaluate the QRC performance.
The scale of QRC is determined by four factors: (i) the number of basis states spanning the Hilbert space, which scales as $N_{c}2^{N_{atom}}$, where $N_{c}$ is the number of involved photon Fock states and $N_{atom}$ is the total number of “measured” and “unmeasured” atoms; (ii) the number of readouts, analogous to the number of neurons in CRC, which is determined by the number of “measured” atoms and related to the index $n$, with a maximum availability of $2N_{atom}+2$; (iii) the number of feedbacks, which is decided by the number of readouts that are used to modify the input through the feedback mechanism, and associated with the index $n_{f}$; and (iv) the implementation of polynomial regression, which incorporates both linear and quadratic terms of the readouts.
@@ II.2 Mackey-Glass Task
The Mackey-Glass task serves as a test for long-term memory, demanding the reservoir to retain past information from the input function to forecast its future behavior. The chaotic nature of the Mackey-Glass time series makes it a challenging problem for machine learning models, forcing them to effectively learn the underlying patterns. The input is generated from the Mackey-Glass equation
@@
where the parameters $\beta=0.2$, $\gamma=0.1$, and $\tau_{M}=17$ are commonly accepted standard values in the chaotic regime [Fujii2017, Dudas2023, Hulser2023]. Following the convention used in Ref. [Hulser2023], we apply a buffer by discarding the first $1000$ time units in Eq. (10). Thus, we consider the discretized function $f_{k}=f\left(t+1000\right)$ as the original input, with discretized time $t=k\Delta t$, as depicted in Fig. 2(a). The time series is sampled with a time step of $\Delta t=1$. The input $f_{k}$ is modified to $\widetilde{f}_{k}$ by the feedback system in each time step, as described by Eq. (1). The target output, $\bar{y}_{k}$, predicting the future of the input function with a time $Delay$, is constructed as $\bar{y}_{k}=f_{k+Delay}$. The readouts of observables illustrated in Fig. 2(c)-(f) show that the reservoir exhibits discernible responses to various input waveforms, which are exemplified by the green shadows corresponding to two different input waveforms. The time series is divided into three intervals for the purposes of memory fading, training and testing. The memory fading phase guarantees that during training and testing, the readouts depend solely on the input function rather than the initial state of the master equation.
@@
Figure 3: Performance testing for the Mackey-Glass task. (a)(b) The tested NRMSE plotted against the number of atoms (or neurons), using all available readouts from the cavity field and all atoms, with no feedback (black), 22 feedbacks via $x_{k1}$ and $x_{k2}$ channels (blue), or 44 feedbacks via $x_{k1}$ to $x_{k4}$ channels (green), where regular linear regression (solid lines) or polynomial regression (dashed lines) is applied. In panel (a), the coupling strength $g_{i}=30$ is fixed for all atoms, and the detunings $\omega_{i}$ vary: $\omega_{i}=20$ for one atom, $\omega_{i}=[0,40]$ for two atoms, $\omega_{i}=[0,20,40]$ for three atoms, $\omega_{i}=[0,10,30,40]$ for four atoms, and $\omega_{i}=[0,10,20,30,40]$ for five atoms. In panel (b), $\omega_{i}=20$ is fixed for all atoms, and $g_{i}$ varies: $g_{i}=30$ for one atom, $g_{i}=[10,50]$ for two atoms, $g_{i}=[10,30,50]$ for three atoms, $g_{i}=[10,20,40,50]$ for four atoms, and $g_{i}=[10,20,30,40,50]$ for five atoms. In panels (a)(b), classical reservoir computing (CRC) (red solid line) uses all available readouts from all neurons in classical echo state networks (ESN) (detailed in the “Supplementary Information”). (c) The tested NRMSE as a function of the number of unmeasured atoms (for QRC) or unmeasured neurons (for CRC), using only the $4$ readout channels (taking 4 measurements) from the cavity field and a particular atom with $g=30$ and $\omega=20$, where the total number of atoms is increased by fixing either $g_{i}$ (solid lines) or $\omega_{i}$ (dashed-dotted lines). (d)(e) Actual (red) and target (blue) outputs from QRC utilizing one atom with no feedback with linear regression, and three atoms with 44 feedbacks with polynomial regression, corresponding to the points marked by letters “d” and “e” in panel (b), respectively. (f)(g) The influence of different selections for the $2$ feedback channels out of the $6$ readout channels in two-atom QRC (black lines), where ($n$,$m$) denotes the feedback channels $x_{kn}$ and $x_{km}$ with $n<m$. The results from one-atom and three-atom QRC using feedback channels $x_{k1}$ and $x_{k2}$ (orange lines) are plotted for reference. Panel (f) fixes $g_{i}$ for all atoms, while panel (g) fixes $\omega_{i}$. (h) Comparison of three methods for training feedback parameters $V_{n_{f}}$: differential evolution, brute force, and brute force plus Nelder Mead, under varying numbers of feedbacks for one-atom QRC with $g_{1}=30$ and $\omega_{1}=20$ using linear regression. Parameters: $Delay=20$, $\kappa=10$, $\omega_{c}=40$, $\epsilon=20$.
@@
The performance enhancement on account of the scalability of QRC is illustrated in Fig. 3. Increasing the number of atoms boosts the performance by expanding both the size of the Hilbert space and the number of available readouts. Figures 3(a)(b) show the results when the number of atoms is increased from $1$ to $5$, utilizing all available readout channels from the cavity field and all the atoms, with $g_{i}$ and $\omega_{i}$ fixed, respectively. This leads to a corresponding increase in the number of readouts (neurons) from $4$ to $12$. The remarkable performance enhancement associated with the feedback mechanism is depicted by the blue ($2$ feedbacks) and green ($4$ feedbacks) lines in Figs. 3(a)(b), while the results from polynomial regression are shown by the dashed lines. A comparison between two extreme cases, one-atom QRC without feedback finished by linear regression and three-atom QRC with 4 feedbacks finished by polynomial regression, is illustrated in Figs. 3(d)(e). This comparison demonstrates a substantial performance improvement resulting from the combination of all the scalable factors.